# UniVideo: Unified Understanding, Generation, and Editing for Videos

### v2

@@ Abstract
Unified multimodal models have shown promising results in multimodal content generation and editing but remain largely limited to the image domain. In this work, we present UniVideo, a versatile framework that extends unified modeling to the video domain. UniVideo adopts a dual-stream design, combining a Multimodal Large Language Model (MLLM) for instruction understanding with a Multimodal DiT (MMDiT) for video generation. This design enables accurate interpretation of complex multimodal instructions while preserving visual consistency. Built on this architecture, UniVideo unifies diverse video generation and editing tasks under a single multimodal instruction paradigm and is jointly trained across them. Extensive experiments demonstrate that UniVideo matches or surpasses state-of-the-art task-specific baselines in text/image-to-video generation, in-context video generation and in-context video editing. Notably, the unified design of UniVideo enables two forms of generalization. First, UniVideo supports task composition, such as combining editing with style transfer, by integrating multiple capabilities within a single instruction. Second, even without explicit training on free-form video editing, UniVideo transfers its editing capability from large-scale image editing data to this setting, handling unseen instructions such as green-screening characters or changing materials within a video. Beyond these core capabilities, UniVideo also supports visual-prompt-based video generation, where the MLLM interprets visual prompts and guides the MMDiT during synthesis. To foster future research, we will release our model and code at https://github.com/KwaiVGI/UniVideo.
Figure 1: UniVideo is a unified system that can understand multi-modal instructions and generate video content. More videos are available on project website.
We are continuing to improve this paper, a refined update will be released in a future version.
@@ 1 Introduction
Despite this progress, unified understanding–generation models remain limited to text and image (Lin et al., 2025; Wu et al., 2025c), leaving video largely underexplored. Existing video generation models primarily address a single text-to-video task and rely on text encoders to process instructions (Wan et al., 2025; Ju et al., 2025; Polyak et al., 2024; Kong et al., 2024), restricting their ability to understand and reason over multimodal instructions (Hu et al., 2024). Meanwhile, video editing methods typically employ task-specific modules or pipelines (Ku et al., 2024; Jiang et al., 2025; Ye et al., 2025b), which makes it difficult to scale across diverse tasks. Consequently, due to the lack of unified modeling, advanced capabilities such as multimodal prompting, in-context video generation, and sophisticated free-form editing remain beyond the reach of any single model.
@@
Based on this unified architecture, we train UniVideo across a wide spectrum of tasks, including text-to-image, text-to-video, image-to-video, in-context video generation, in-context video editing, and image editing. As a unified system, UniVideo not only understands multimodal instructions and distinguishes between tasks but also achieves improvements over state-of-the-art task-specific methods. Thanks to unified training, UniVideo generalizes to novel task compositions unseen during training, such as deleting one identity while swapping another within a single instruction. More importantly, although UniVideo is not trained on free-form video editing data, it demonstrates generalization ability transfer from image editing to free-form video editing (e.g., change material and weather), highlighting the effectiveness of our unified video understanding and generation framework.
Furthermore, UniVideo retains the strong visual understanding capability of its underlying frozen MLLM. By leveraging the MLLM’s autoregressive reasoning and language generation abilities, UniVideo can effectively interpret ambiguous and complex multimodal instructions that require joint vision–language understanding, such as turning visual prompting into in-context video generation tasks. Since its text generation ability originates from a frozen MLLM, UniVideo should be regarded as an assembled multimodal generative system capable of producing images, videos, and text, rather than a native multimodal model trained end-to-end (Ma et al., 2025b; Deng et al., 2025).
@@
We introduce UniVideo, a powerful multimodal generative model that unifies understanding, generation, and editing of videos within a single framework. To build UniVideo, we propose a dual-stream architecture that combines the multimodal reasoning capabilities of the MLLM with the generation strengths of the MMDiT. Unlike prior task-specific or modality-restricted approaches, UniVideo can interpret multimodal instructions, distinguish between diverse tasks, and achieve state-of-the-art performance across a wide range of benchmarks.
2) We demonstrate that UniVideo generalizes to unseen tasks and novel task compositions without ad hoc designs, highlighting the benefits of a unified framework.
@@ 2.1 Model Architecture
As demonstrated in Figure 2, UniVideo consists of two main components: a multimodal large language model (MLLM) and a multimodal DiT (MM-DiT). The MLLM handles visual–textual understanding, taking text, image, and video inputs and producing text responses. The MM-DiT focuses on visual generation with two branches: one incorporates high-level semantic information from the MLLM, while the other integrates fine-grained reconstruction signals from a VAE. Specifically, we extract the last-layer hidden states of the MLLM, which encode rich semantic features of the multimodal input. These are aligned to the input space of the MM-DiT via a trainable connector and fed into its understanding stream. In parallel, visual signals are encoded by the VAE and passed into the MM-DiT generation stream to preserve fine details. This design enables strong semantic grounding together with high-fidelity visual detail, which is especially important for video editing and identity-preserving in-context generation.
@@ 2.2 Unifying Multiple Tasks
We standardize multimodal instructions by assigning each visual input an ID tag, as illustrated in Figure 1. For text-to-video (T2V), the text input is processed by the MLLM, while the noisy video is fed into the MM-DiT. For image-to-video (I2V), both the image and text are processed by the MLLM, whereas the image and noisy video are provided to the MM-DiT. For in-context video generation (MultiID2V) and in-context video editing (ID-V2V), multiple visual conditions are often available, such as several reference images together with a reference video. Each visual signal is encoded with the VAE, padded to a uniform shape, concatenated along the temporal axis, and then processed with self-attention. Unlike prior approaches that introduce task-specific bias embeddings (Ye et al., 2025b) or context adapter modules (Jiang et al., 2025), we avoid task-specific customization. To help the MM-DiT distinguish between condition latents and noisy video latents, we apply 3D positional embeddings, which preserve the spatial indices across frames while incrementing only the temporal dimension. In practice, we find this strategy more effective than Qwen2-VL’s MRoPE (Wang et al., 2024a), which offsets all axes whenever a new visual input is introduced.
@@
2.3 Understanding Visual Prompt
2.4 Training Strategy
Stage 1. Connector alignment between MLLM and MMDiT. In this stage, we train only the MLP connector while keeping both the MLLM and MMDiT frozen. Training is performed on $\mathcal{O}(40)\text{M}$ pretraining samples across text-to-image (T2I) and $\mathcal{O}(10)\text{M}$ text-to-video (T2V) generation tasks, as well as an image-reconstruction task in which only images from the text-to-image dataset are fed into the MLLM and the MMDiT reconstructs the image using visual features from the MLLM. After this stage, UniVideo can generate images and videos conditioned on text or image inputs from the MLLM.
Stage 2. Fine-tuning MMDiT on T2I and T2V. In this stage, we keep the MLLM frozen and fine-tune the connector and MMDiT on $\mathcal{O}(10)\text{K}$ high-quality T2I and T2V samples. After this stage, UniVideo achieves performance comparable to the MMDiT backbone that uses its own text encoder.
Stage 3. Multi-task training. Finally, we extend training to include in-context generation (multi-ID-to-video), in-context video editing, image editing and image-to-video tasks, alongside the previous T2I and T2V tasks. We keep the MLLM frozen and only train the connector and MMDiT. This stage enables UniVideo to unify a broad range of video generation and editing tasks under multimodal instruction. Details of task decomposition, training setting and dataset construction are provided in Table 1 and Table 2.
Table 1: Overview of the multimodal training data used for UniVideo. Each task is characterized by its input modalities and number of examples.
Task Input #Examples
Text to Image txt $\mathcal{O}(40)\text{M}$
Text to Image (High Quality) txt $\mathcal{O}(10)\text{K}$
Image Reconstruction image $\mathcal{O}(40)\text{M}$
Text to Video txt $\mathcal{O}(10)\text{M}$
Text to Video (High Quality) txt $\mathcal{O}(10)\text{K}$
Image to Video img+txt $\mathcal{O}(10)\text{K}$
Image Editing img+txt $\mathcal{O}(1)\text{M}$
Image Style Transfer img+txt $\mathcal{O}(10)\text{K}$
In-Context Video Editing (swap, addition, delete, style) ref-img $\times$ n + video + txt $\mathcal{O}(10)\text{K}$
In-Context Video Generation ref-img $\times$ n + txt $\mathcal{O}(10)\text{K}$
In-Context Image Style Transfer ref-img $\times$ n + img + txt $\mathcal{O}(10)\text{K}$
@@ 2.4 Training Strategy
# Training samples $\mathcal{O}(50)\text{M}$ $\mathcal{O}(10)\text{K}$ Mixed tasks (Table 1)
@@
Data sampling ratio
Text to Image 0.7 0.0 0.0
Text to Image(High Quality) 0.0 0.7 0.05
Text to Video 0.2 0.0 0.0
Text to Video(High Quality) 0.0 0.2 0.05
Image Reconstruction 0.1 0.1 0.0
Image to Video 0.0 0.0 0.1
Image Editing 0.0 0.0 0.3
Image Style Transfer 0.0 0.0 0.1
In-Context Video Editing 0.0 0.0 0.1
In-Context Video Generation 0.0 0.0 0.2
In-Context Image Style Transfer 0.0 0.0 0.1
@@ 3 Experiments
In this section, we first describe the implementation details in subsection 3.1. Then, we present main results in subsection 3.2. We conduct a comprehensive benchmark of UniVideo with SoTA methods across a broad spectrum of video understanding and generation tasks. Our results show that UniVideo’s strong unified capabilities across all settings. Next, we demonstrate the zero shot generalization ability of UniVideo and analysis the visual prompt understanding ability in subsection 3.3. Finally, we validate the design choices of UniVideo through ablation studies in subsection 3.4.
@@ 3.1 Implementation Details
We adopt qwen2.5VL-7B (Bai et al., 2025) as the MLLM backbone and HunyuanVideo-T2V-13B (Kong et al., 2024) as the MMDiT backbone. The original HunyuanVideo use two text encoders; we remove them and instead use qwen2.5VL as the unified multi-modal embedder. To align feature dimensions between qwen2.5VL and HunyuanVideo, we apply an MLP with a $4\times$ expansion. Training is conducted on 32 H100 GPUs.
@@ 3.2.1 Visual Understanding and Generation
Emu3 58.5 31.6 37.2 80.96
Show-o2 79.3 48.9 56.6 81.34
@@ 3.2.2 In-Context Video Generation
Benchmark: Following FullDiT (Ju et al., 2025) and OmniGen2 (Wu et al., 2025c), we construct a test set covering both single-ID and multi-ID video generation scenarios. In the single-ID setting, a subject may have multiple reference images (e.g., different viewpoints of a person or object). In the multi-ID setting, the references include 2–4 distinct identities. Details are provided in the Appendix.
Metrics: We conduct both human evaluations and automatic metric assessments. For human evaluation, we follow the protocols of Instruct-Imagen (Hu et al., 2024) and OmniGen2 (Wu et al., 2025c) to perform a systematic study. Each sample is rated by at least three annotators on (i) subject consistency (SC), (ii) prompt following (PF), and (iii) overall video quality (Overall). Scores in each category are drawn from $\{0,0.5,1\}$, where 0 indicates inconsistency or extremely poor quality, and 1 indicates full consistency or high quality. For automatic evaluation, we adopt three metrics from VBench (Huang et al., 2024): smoothness, dynamics, and aesthetics.
@@
Table 4: Quantitative comparison on In-Context Generation. Human evaluation includes Subject Consistency (SC), Prompt Following (PF), and Overall scores. Automatic metrics measure video quality in terms of Smoothness, Dynamics, and Aesthetics. Best results are shown in bold, and second-best are underlined. UniVideo achieves superior or competitive performance across all metrics compared to the SoTA methods and commercial models and in particular be the best for SC.
@@
VACE 0.31 0.65 0.42 0.922 40.341 5.426
Kling1.6 0.68 0.95 0.88 0.938 86.641 5.896
Pika2.2 0.45 0.43 0.15 0.928 104.768 5.125
@@
VACE 0.48 0.53 0.48 0.53 65.606 5.941
Kling1.6 0.73 0.45 0.95 0.916 61.856 6.034
Pika2.2 0.71 0.48 0.43 0.898 76.796 5.176
@@
VACE 0.513 0.105 0.103 0.947 51.343 5.693
UNIC 0.598 0.245 0.216 0.961 11.070 5.627
Kling1.6 0.632 0.287 0.246 0.993 1.025 5.798
Pika2.2 0.692 0.399 0.253 0.951 261.443 5.591
@@
VACE 0.703 0.391 0.218 0.960 29.001 5.961
UNIC 0.725 0.429 0.242 0.971 7.500 6.056
Kling1.6 0.707 0.437 0.211 0.995 0.518 6.042
Pika2.2 0.704 0.406 0.211 0.967 30.812 5.097
AnyV2V 0.605 0.229 0.218 0.917 7.596 4.842
@@
VACE 20.601 0.874 0.206 0.968 16.146 5.637
UNIC 19.171 0.817 0.217 0.970 10.934 5.493
Kling1.6 15.476 0.888 0.208 0.998 0.663 4.965
AnyV2V 19.504 0.869 0.205 0.964 4.980 5.325
VideoPainter 22.987 0.920 0.212 0.957 13.759 5.403
@@
AnyV2V 0.207 43.299 0.195 0.937 9.227 4.640
StyleMaster 0.306 38.213 0.188 0.952 9.758 5.121
UNIC 0.197 36.198 0.215 0.932 11.569 5.045

### v3

@@ Abstract
Unified multimodal models have shown promising results in multimodal content understanding and generation but remain largely limited to the image domain. In this work, we present UniVideo, a versatile framework that extends unified modeling to the video domain. UniVideo adopts a dual-stream design, combining a Multimodal Large Language Model (MLLM) for visual understanding with a Multimodal DiT (MMDiT) for visual generation. This design preserves the MLLM’s original text generation capabilities, enables accurate interpretation of complex multimodal instructions, and maintains visual consistency in the generated content. Built on this architecture, UniVideo unifies diverse video generation and editing tasks under a single multimodal instruction paradigm and is jointly trained across them. Extensive experiments demonstrate that UniVideo matches or surpasses state-of-the-art task-specific baselines in visual understanding, text/image-to-video generation, in-context video generation and in-context video editing. Notably, the unified design of UniVideo enables two forms of generalization. First, UniVideo supports task composition, such as combining editing with style transfer, by integrating multiple capabilities within a single instruction. Second, even without explicit training on free-form video editing, UniVideo transfers its editing capability from large-scale image editing data to this setting, handling unseen instructions, such as changing the environment or altering materials within a video. Beyond these core capabilities, UniVideo also supports visual-prompt-based video generation, where the MLLM interprets visual prompts and guides the MMDiT during synthesis. To foster future research, we released our model and code at https://github.com/KlingTeam/UniVideo.
Figure 1: UniVideo is a unified system that can understand multi-modal instructions and generate multi-modal content. More videos are available on project website.
@@ 1 Introduction
Despite this progress, unified understanding–generation models remain limited to text and image (Lin et al., 2025; Wu et al., 2025c), leaving video largely underexplored. Existing video generation models primarily address a single text-to-video task and rely on text encoders to process instructions (Wan et al., 2025; Ju et al., 2025b; Polyak et al., 2024; Kong et al., 2024), restricting their ability to understand and reason over multimodal instructions (Hu et al., 2024). Meanwhile, video editing methods typically employ task-specific modules or pipelines (Ku et al., 2024; Jiang et al., 2025; Ye et al., 2025b), which makes it difficult to scale across diverse tasks. Consequently, due to the lack of unified modeling, advanced capabilities such as multimodal prompting, in-context video generation, and sophisticated free-form editing remain beyond the reach of any single model.
@@
Based on this unified architecture, we train UniVideo across a wide spectrum of tasks, including text-to-image, text-to-video, image-to-video, in-context video generation, in-context video editing, and image editing. As a unified system, UniVideo not only understands multimodal instructions and distinguishes between tasks but also achieves improvements over state-of-the-art task-specific methods. Thanks to unified training, UniVideo generalizes to novel task compositions unseen during training, such as deleting one identity while swapping another within a single instruction. More importantly, although UniVideo is not trained on free-form video editing data, it demonstrates generalization ability transfer from image editing to free-form video editing (e.g., changing object materials or modifying weather conditions), highlighting the effectiveness of our unified video understanding and generation framework.
Furthermore, UniVideo retains the strong visual understanding and text generation capability of its underlying frozen MLLM. By leveraging the MLLM’s autoregressive reasoning and language generation abilities, UniVideo can effectively interpret ambiguous and complex multimodal instructions that require joint vision–language understanding, such as turning visual prompting inputs into in-context video generation tasks or image-to-video generation tasks. Since UniVideo’s text generation ability originates from a frozen MLLM, UniVideo should be regarded as a post-trained unified multimodal generative system(Wu et al., 2025c; Pan et al., 2025) capable of producing images, videos, and text, rather than a unified model trained from scratch(Ma et al., 2025b; Deng et al., 2025).
@@
We introduce UniVideo, a multimodal generative model that unifies understanding, generation, and editing of videos within a single framework. To build UniVideo, we propose a dual-stream architecture that combines the multimodal reasoning capabilities of the MLLM with the generation strengths of the MMDiT. Unlike prior task-specific or modality-restricted approaches, UniVideo can interpret multimodal instructions, distinguish between diverse tasks, and achieve state-of-the-art performance across a wide range of benchmarks.
2) We systematically study the key design choices that enable this unified framework, including connector architectures, generator designs, and multimodal conditioning strategies, and provide empirical evidence for their effectiveness.
3) We demonstrate that UniVideo generalizes to unseen tasks and novel task compositions without ad hoc designs, highlighting the benefits of a unified framework.
@@ 2.1 Model Architecture
As demonstrated in Figure 2, UniVideo consists of two main components: a multimodal large language model (MLLM) and a multimodal DiT (MM-DiT). The MLLM handles visual–textual understanding, taking text, image, and video inputs and optionally producing text responses. The MM-DiT focuses on visual generation with two branches: one incorporates high-level semantic information from the MLLM, while the other integrates fine-grained reconstruction signals from a VAE. Specifically, we extract the last-layer hidden states of the MLLM, which encode rich semantic features of the multimodal input. These are aligned to the input space of the MM-DiT via a trainable connector and fed into its understanding stream. In parallel, visual signals are encoded by the VAE and passed into the MM-DiT generation stream to preserve fine details. This design enables strong semantic grounding together with high-fidelity visual detail, which is especially important for video editing and identity-preserving generation. We provide a model design analysis in subsection 3.2.
@@ 2.2 Unifying Multiple Tasks
We unify diverse multimodal tasks through natural language instructions, as illustrated in Figure 1. For text-to-video (T2V), the text input is processed by the MLLM, while the noisy video is fed into the MM-DiT. For image-to-video (I2V), both the image and text are processed by the MLLM, whereas the image and noisy video are provided to the MM-DiT. For in-context video generation (MultiID2V) and in-context video editing (ID-V2V), multiple visual conditions are often available, such as several reference images together with a reference video. Each visual signal is encoded with the VAE, padded to a uniform shape, concatenated along the temporal axis, and then processed with self-attention. Unlike prior approaches that introduce task-specific bias embeddings (Ye et al., 2025b) or context adapter modules (Jiang et al., 2025), we avoid task-specific customization. To help the MM-DiT distinguish between condition latents and noisy video latents, we apply 3D positional embeddings, which preserve the spatial indices across frames while incrementing only the temporal dimension. In practice, we find this strategy more effective than Qwen2-VL’s MRoPE (Wang et al., 2024a), which offsets all axes whenever a new visual input is introduced.
@@
2.3 Understanding Visual Prompt
2.4 Training Strategy
Stage 1. Connector alignment between MLLM and MMDiT. In this stage, we train only the MLP connector while keeping both the MLLM and MMDiT frozen. Training is performed on pretraining samples across text-to-image (T2I) and text-to-video (T2V) generation tasks, as well as an image-reconstruction task in which only images from the text-to-image dataset are fed into the MLLM and the MMDiT reconstructs the image using visual features from the MLLM. After this stage, UniVideo can generate images and videos conditioned on text or image inputs from the MLLM.
Stage 2. Fine-tuning MMDiT on T2I and T2V. In this stage, we keep the MLLM frozen and fine-tune the connector and MMDiT on small scale high-quality T2I and T2V samples. After this stage, UniVideo achieves performance comparable to the MMDiT backbone that uses its own text encoder.
Stage 3. Multi-task Training. Finally, we extend training to include in-context generation (multi-ID-to-video), in-context video editing (modifying the input video based on a reference image, such as ID swapping, ID addition, ID deletion, or style transfer), image editing and image-to-video tasks, alongside the previous text-to-image (T2I) and text-to-video (T2V) tasks. We keep the MLLM frozen and only train the connector and MMDiT. This stage enables UniVideo to unify a broad range of video generation and editing tasks under multimodal instruction. Details of training setting is provided in Table 1.
Figure 4: Three design choices for aligning the MLLM with the diffusion generator in Stage 1 training. We keep the MLLM fixed and vary the connector and DiT architecture across three variants: (a) the DiT uses cross-attention for text conditioning, where we replace its original text encoder with an MLP layer that aligns the final hidden states from the MLLM; (b) building upon (a), we introduce a learnable query design and extract the final hidden states from these learnable queries; and (c) our UniVideo architecture employs an MMDiT design that leverages self-attention for text conditioning.
Figure 5: Qualitative comparison of design choices for aligning the MLLM with the diffusion generator in Stage 1 training. In all settings, the MLLM is kept frozen. (a) Cross-Attention DiT: we train the MLP connector and DiT; (b.1) Cross-Attention DiT with Learnable Query: following (Pan et al., 2025), we train the learnable query tokens, MLP connector, and DiT; (b.2) similar to (b.1), but the DiT is frozen while only the learnable query tokens and MLP connector are trained; (c) UniVideo (MMDiT): only the MLP connector is trained, with all other components frozen. All variants are trained for 15K steps. Among all variants, UniVideo (MMDiT) demonstrates the best prompt alignment.
Figure 6: Two design variants for multimodal conditioning in Stage 3 multi-task training. We study two multimodal conditioning strategies: (a) extracting the final hidden states of image, video, and text tokens produced by the MLLM; and (b) adopting a learnable query design and using the final hidden states of the learnable queries together with the final hidden states of the text tokens.
@@ 3 Experiments
In this section, we first describe the implementation details of UniVideo in subsection 3.1. Next, we discuss the design choices for aligning the MLLM and the Diffusion generator in subsection 3.2. Then, We present the main results in subsection 3.3. We conduct a comprehensive benchmark of UniVideo with SoTA methods across a broad spectrum of video understanding and generation tasks. Our results show that UniVideo’s strong unified capabilities across all settings. Next, we demonstrate the zero shot generalization ability of UniVideo and analysis the visual prompt understanding ability in subsection 3.4. Finally, we validate the design choices of UniVideo through ablation studies in subsection 3.5.
@@ 3.1 Implementation Details
We adopt qwen2.5VL-7B (Bai et al., 2025) as the MLLM backbone and HunyuanVideo-T2V-13B (Kong et al., 2024) as the MMDiT backbone. The original HunyuanVideo use two text encoders; we remove them and instead use qwen2.5VL as the unified multi-modal embedder. To align feature dimensions between qwen2.5VL and HunyuanVideo, we apply an MLP with a $4\times$ expansion.
3.2 Model Design
Our model design study addresses the following question: What is the most effective approach for aligning a pretrained MLLM with a diffusion generator during Stage 1 training?
We investigate three design choices for aligning the pretrained MLLM with the diffusion generator in Stage 1. Throughout this stage, the MLLM remains frozen, while we vary the connector and DiT architectures across three variants as shown in  Figure 4.
(a) Cross-attention DiT. The first variant adopts a cross-attention–based DiT for text conditioning, where we replace its original text encoder with an MLP connector that projects the final hidden states from the MLLM into the DiT text embedding space. Both the MLP and DiT are trained.
(b) Cross-attention DiT with Learnable query. Building upon (a), we use a learnable query mechanism following Pan et al. (2025). Specifically, we extract the final hidden states of learnable queries from the MLLM, which are then passed through an MLP layer and used to replace the original text conditioning in the DiT’s cross-attention module. We test two variants: (1) jointly training the learnable queries, MLP layer, and DiT (as in Pan et al. (2025)); and (2) training only the learnable queries and MLP while keeping the DiT frozen.
(c) UniVideo architecture. The main difference in this variant lies in its use of MMDiT, which employs self-attention for joint text–video interaction instead of cross-attention. We replace MMDiT’s original text encoder with an MLP connector that projects the final hidden states from the MLLM into the MMDiT’s text embedding space. Only the MLP layer is trained, while both the MLLM and MMDiT remain frozen.
For the cross-attention variants, we use an internal model with an architecture similar to (Wan et al., 2025), originally based on a T5 text encoder(Raffel et al., 2020), which we replace with Qwen2.5-VL. For UniVideo, we follow the implementation details described in subsection 3.1. All variants are trained for 15K steps, and the qualitative results are presented in Figure 5.
Our findings show that the cross-attention variants require unfreezing the DiT generator to achieve effective alignment with the MLLM, as evidenced by the comparison between (b.2) and (b.1). Nevertheless, even after unfreezing, variants (a) and (b.1) exhibit limited text-following ability—particularly for compositional object prompts. In contrast, the UniVideo architecture achieves efficient and robust alignment by training only the MLP connector.
We study two UniVideo variants in Stage 3 training. The MLLM is kept frozen, while we vary the connector design across two variants, as illustrated in Figure 6.
(a) UniVideo (hidden). In this variant, we extract the final-layer hidden states of all image, video, and text tokens produced by the MLLM. These token representations are used as inputs to the understanding branch of MMDiT. During training, only the MLP connector and the MMDiT are updated, while the MLLM remains frozen.
(b) UniVideo (query). This variant adopts a learnable query mechanism following Pan et al. (2025). Specifically, we extract the final hidden states of the learnable query tokens from the MLLM, together with the final hidden states of the text tokens. In this setting, we train the learnable queries, the MLP connector, and the MMDiT, while keeping the MLLM frozen.
UniVideo (query) can be more computationally efficient at inference time due to its fixed number of query tokens. By default, we use 512 learnable queries, which reduces the number of conditioning tokens compared to using all the multimodal hiddens from the MLLM. This efficiency gain is particularly beneficial for tasks where video inputs dominate the MMDiT conditioning, such as video editing.
During training, however, the UniVideo (query) variant requires backpropagation through the MLLM computation graph in order to optimize the learnable queries, which incurs additional memory overhead. In contrast, the UniVideo (hidden) variant does not require gradient flow through the MLLM and is therefore more memory-efficient during training.
Unless otherwise specified, we report results using the UniVideo (hidden) variant throughout this paper.
@@ 3.3.1 Visual Understanding and Generation
Emu3 (Wang et al., 2024b) 58.5 31.6 37.2 80.96
TokenFlow-XL (Qu et al., 2025) 76.8 43.2 48.2 $\times$
Janus (Wu et al., 2025b) 69.4 30.5 34.3 $\times$
JanusFlow (Ma et al., 2025b) 74.9 29.3 30.9 $\times$
OmniGen2 (Wu et al., 2025c) 79.1 53.1 61.8 $\times$
Show-o (Xie et al., 2024) - 26.7 - $\times$
BAGEL (Deng et al., 2025) 85.0 55.3 67.2 $\times$
Show-o2 (Xie et al., 2025) 79.3 48.9 56.6 81.34
@@ 3.3.2 In-Context Video Generation
Benchmark: Following FullDiT (Ju et al., 2025b) and OmniGen2 (Wu et al., 2025c), we construct a test set covering both single-ID and multi-ID video generation scenarios. In the single-ID setting, a subject may have multiple reference images (e.g., different viewpoints of a person or object). In the multi-ID setting, the references include 2–4 distinct identities. Details are provided in the Appendix.
Metrics: We conduct both human evaluations and automatic metric assessments. For human evaluation, we follow the protocols of Instruct-Imagen (Hu et al., 2024) and OmniGen2 (Wu et al., 2025c) to perform a systematic study. Each sample is rated by at least three annotators on (i) subject consistency (SC), (ii) prompt following (PF), and (iii) overall video quality (VQ). Scores in each category are drawn from $\{0,0.5,1\}$, where 0 indicates inconsistency or extremely poor quality, and 1 indicates full consistency or high quality. For automatic evaluation, we adopt three metrics from VBench (Huang et al., 2024): smoothness, and aesthetics.
@@
Table 3: Quantitative comparison on In-Context Generation. Human evaluation includes Subject Consistency (SC), Prompt Following (PF), and Overall Video Quality (VQ). Automatic metrics measure video quality in terms of Smoothness and Aesthetics. Best results are shown in bold, and second-best are underlined. UniVideo achieves superior or competitive performance across all metrics compared to the SoTA methods and commercial models and in particular be the best for SC.
@@
VACE 0.31 0.65 0.42 0.922 5.426
Kling1.6 0.68 0.95 0.88 0.938 5.896
Pika2.2 0.45 0.43 0.15 0.928 5.125
@@
VACE 0.48 0.53 0.48 0.53 5.941
Kling1.6 0.73 0.45 0.95 0.916 6.034
Pika2.2 0.71 0.48 0.43 0.898 5.176
@@
VACE 0.513 0.105 0.103 0.947 5.693
UNIC 0.598 0.245 0.216 0.961 5.627
Kling1.6 0.632 0.287 0.246 0.993 5.798
Pika2.2 0.692 0.399 0.253 0.951 5.591
@@
VACE 0.703 0.391 0.218 0.960 5.961
UNIC 0.725 0.429 0.242 0.971 6.056
Kling1.6 0.707 0.437 0.211 0.995 6.042
Pika2.2 0.704 0.406 0.211 0.967 5.097
AnyV2V 0.605 0.229 0.218 0.917 4.842
@@
VACE 20.601 0.874 0.206 0.968 5.637
UNIC 19.171 0.817 0.217 0.970 5.493
Kling1.6 15.476 0.888 0.208 0.998 4.965
AnyV2V 19.504 0.869 0.205 0.964 5.325
VideoPainter 22.987 0.920 0.212 0.957 5.403
@@
AnyV2V 0.207 43.299 0.195 0.937 4.640
StyleMaster 0.306 38.213 0.188 0.952 5.121
UNIC 0.197 36.198 0.215 0.932 5.045